# ZS-VCOS: Zero-Shot Outperforms Supervised Video Camouflaged Object Segmentation

### v1

@@ Abstract
Camouflaged object segmentation presents unique challenges compared to traditional segmentation tasks, primarily due to the high similarity in patterns and colors between camouflaged objects and their backgrounds. Effective solutions to this problem have significant implications in critical areas such as pest control, defect detection, and lesion segmentation in medical imaging. Prior research has predominantly emphasized supervised or unsupervised pre-training methods, leaving zero-shot approaches significantly underdeveloped. Existing zero-shot techniques commonly utilize the Segment Anything Model (SAM) in automatic mode or rely on vision-language models to generate cues for segmentation; however, their performances remain unsatisfactory, likely due to the similarity of the camouflaged object and the background. Optical flow, commonly utilized for detecting moving objects, has demonstrated effectiveness even with camouflaged entities. Our method integrates optical flow, a vision-language model, and SAM 2 into a sequential pipeline. Evaluated on the MoCA-Mask dataset, our approach achieves outstanding performance improvements, significantly outperforming existing zero-shot methods by raising the F-measure ($F_{\beta}^{w}$) from 0.296 to 0.628. Remarkably, our approach also surpasses supervised methods, increasing the F-measure from 0.476 to 0.628. Additionally, evaluation on the MoCA-Filter dataset demonstrates an increase in the success rate from 0.628 to 0.697 when compared with FlowSAM, a supervised transfer method. A thorough ablation study further validates the individual contributions of each component.  [^1] More details can be found on https://github.com/weathon/vcos.
@@ 1 Introduction
Figure 1: Evolution of $F_{\beta}^{w}$ scores over time for supervised and zero-shot methods on an animal dataset. The $F_{\beta}^{w}$ metric is selected as the most representative and consistently reported across all compared methods. Most zero-shot approaches utilize prior knowledge by explicitly instructing models to detect animals, except for CVP and SAM-2L Auto. Our zero-shot method notably surpasses all previous zero-shot approaches and even outperforms supervised methods, achieving performance close to the test-time semi-supervised upper bound.
Camouflaged object detection and segmentation (COD and COS) pose significant challenges beyond traditional object detection and segmentation tasks. This increased difficulty primarily stems from the inherent nature of camouflaged objects, which are visually similar in patterns, colors, and textures to their backgrounds [9] [50], effectively blending into their surroundings and complicating accurate identification and delineation. Despite these challenges, effective solutions for COD and COS have considerable real-world significance, especially in critical fields such as defect detection [19], pest control [38], and medical imaging for lesion segmentation [13].
@@
However, camouflage-related tasks in both static and dynamic contexts remain relatively underdeveloped compared to traditional detection and segmentation methods. Most prior work in this area has focused on supervised learning, relying on complex architectures and labeled data. Yet, even these supervised models often struggle with camouflaged objects due to the lack of distinct features. On the other hand, zero-shot methods, which avoid training by using large pre-trained models like SAM and vision-language models, are severely less explored and currently perform worse than supervised methods.
To address this gap, we propose a method that integrates optical flow, a vision-language model, and SAM in a modular pipeline. Each stage of the pipeline uses the output of the previous one to refine its segmentation cues. Rather than relying on any training or fine-tuning, our approach operates in a zero-shot setting and achieves strong performance. On the MoCA-Mask dataset, our method improves mIoU from 0.273 (baseline zero-shot methods) to 0.561. It also outperforms multiple supervised methods, which typically score around 0.422. Furthermore, on the MoCA-Filtered dataset, our method raises the detection success rate from 0.628 to 0.697. These gains highlight the effectiveness of combining motion-based cues with strong foundational models.
In summary, our contributions are (1) a zero-shot framework for camouflaged object segmentation in video that surpasses supervised baselines, (2) extensive experimentation on different components and prompting strgies of our methods, and (3) insights demonstrating that properly designed zero-shot pipelines can not only compete with but in some cases outperform traditional supervised approaches.
Figure 2: FLow Chart of Our Method
Figure 3: Visual Comparison Of Our Methods and Pervious Supervised Methods
@@ 2.1 Optical Flow
Optical flow is a technique used to measure pixel movement in videos. It has been used in video processing or recognition for a long time; one of the most significant works is the two-stream network published in 2014 [40]. There are two types of optical flow: sparse and dense optical flow. Sparse optical flow gives a movement vector for points of interest in the image, whereas dense optical flow estimates movement for all pixels in the image. One of the most famous sparse optical flows is the Lucas–Kanade method [26], which uses the assumption that local pixels have similar motion. It can be used in camera motion estimation for panoramic image generation or motion compensation. Dense optical methods, like RAFT [43] and GMFlow [53], can provide movement information for every pixel in the frame, and it has demonstrated promising performance in camouflaged object detection based on movement differences between foreground and background.
Method Pub. Setting $S_{\alpha}$ $F_{\beta}^{w}$ MAE
SLT-Net [9] CVPR 22 SV Tr 0.656 0.357 0.021
ZoomNeXt [35] TPAMI 24 SV Tr 0.734 0.476 0.010
TSP-SAM(M+B) [16] CVPR 24 SV Tr 0.689 0.444 0.008
Gao et. al [14] arXiv 25 SV Tr 0.709 0.451 0.008
SAM2 Tracking [41] arXiv 24 SV Te 0.804 0.691 0.004
SAM-PM [29] CVPRW 24 SV Tr+Te 0.728 0.567 0.009
Finetuned SAM2-T + Prompts [58] arXiv 24 SV Tr+Te 0.832 0.726 0.005
CVP [42] ACM MM 24 ZS 0.569 0.196 0.031
SAM-2-L Auto [58] arXiv 24 ZS 0.447 0.198 0.250
LLaVA + SAM2-L [58] arXiv 24 ZS w/ PK 0.622 0.296 0.047
Shikra + SAM2-L [58] arXiv 24 ZS w/ PK 0.495 0.132 0.107
Ours - ZS w/ PK 0.776 0.628 0.008
Table 1: Performance comparison on the MoCA-Mask dataset [9]. ”SV Tr” denotes supervised training, and ”SV Te” denotes supervised testing, where one frame from the video was provided to the model along with prompts. ”ZS” indicates zero-shot learning, while ZS w/ PK means zero-shot with prior knowledge (since the model already knows it is looking for animals). Our method significantly outperforms all zero-shot and even supervised methods.
@@ 2.2.1 Methods Leveraging Optical Flow as Motion Cues
Brox et al. [4] utilized long-term optical flow trajectories combined with clustering for segmenting videos. Ochs et al. [33] applied flow-based motion cues to resolve ambiguities in color-based segmentation. Xiao et al. [51] employed optical flow cues to reinforce target frame representations. Yang et al. [57] uses both optical flow and RGB input to assist video object segmentation. FlowI-SAM and FlowP-SAM [52] utilized optical flow either exclusively as input (FlowI-SAM) or as a prompt guiding segmentation of RGB frames (FlowP-SAM). Both FlowI-SAM and FlowP-SAM handle standard and camouflaged objects effectively.
@@ 2.4 Camouflage Object Detection and Segmentation
Unlike regular object detection and segmentation, camouflage object tasks are significantly more challenging because the foreground usually seamlessly blends into the background. There are two tasks in camouflage object detection: image-based camouflage object segmentation (usually referred to as COS) and video-based camouflage object segmentation (VCOS). They could also be extended to segmentation, namely COS and VCOS. VCOD/S allows the model to use motion cues to detect the foreground but also brings in the challenges of temporal changes [50].
@@ 2.4.1 Datasets
There are two major datasets and 3 variants in video camouflage object detection: Camouflaged Animal Dataset (CAD) [3] and Moving Camouflaged Animal Dataset (MoCA) [21]. However, MoCA is an object detection dataset and not a segmentation dataset, and it contains some non-camouflaged animals or animals that do not have locomotion. Thus, two variances of MoCA are proposed. MoCA-filtered [56] and MoCA-Mask [9].
MoCA-filtered mainly removed non-locomotive videos from the dataset, with additional processing such as cropping away logos and borders, resampling frames, and incorporating the bounding boxes. Since it still lacks segmentation masks, papers using this dataset ([56] [52]) use the detection success rate based on the IoU threshold to evaluate the results. MoCA-Mask improves MoCA by removing scenes with obvious animals and converting bounding boxes into masks. In addition to ground truth masks every 5 frames, they also used bidirectional optical flow to generate pseudo masks for unlabeled frames.
@@ 2.4.2 Algorithms
SLT-Net [9] is a supervised trained and unsupervised tested model. It argued that when using optical flow and homography, the error might be accumulated from both the motion estimation and segmentation. Thus, they proposed to use a unified framework for both motion estimation and segmentation. Additionally, they used a long-term spatio-temporal transformer to refine short-term predictions, although this long-term module provides marginal improvement. ZoomNeXt [35] is an improved version of ZoomNet [34], mainly adapted from image-based COS to video-based COS and improved performance by introducing more structural extensions. ZoomNet and ZoomNeXt both use zooming to capture features at different scales. Similar to SLT-Net, they are both supervised trained and unsupervised test methods.
Previous studies occasionally failed to differentiate clearly between test-time semi-supervised and test-time unsupervised tasks despite their distinct levels of challenge. For example, SAM-PM [29], requiring supervision during both training and inference, reported state-of-the-art results compared with SLT-Net [9]. However, SLT-Net operates under supervised training but unsupervised testing conditions. This fundamental difference in evaluation criteria renders direct comparisons between these two methods somewhat inequitable.
@@
For zero-shot unsupervised testing (neither training nor inference supervision), Chain of Vision Perception (CVP) [42] represents an early effort employing vision-language models (VLMs) for COD/S tasks, with a primary focus on images rather than videos. CVP prompts a vision-language model to identify the location of camouflaged objects. Subsequently, these locations are refined and given to a segmentation model. Properly designed prompting can further enhance the model’s performance. CVP achieved higher performance than several supervised methods on datasets such as CAMO [22], COD10K [12], and NC4K [27]. However, its results on the MoCA-Mask were suboptimal, with a weighted F-score ($F_{\beta}^{w}$) of 0.196. Zhou et al. [58] improved upon this by employing LLaVA [24] or Shikra [5] as the vision-language model and utilizing SAM-2 for segmentation. A similar approach is evident in Grounded SAM [25], which integrates Grounding DINO and SAM for open-vocabulary segmentation of regular objects.
@@
Motion Detection Mean Subtraction Momentum Tracking $S_{\alpha}$ $E_{\phi}$ mIoU
@@ 3.1 Motion Detection
Our method builds upon LangGas [15]. Previous studies, including LangGas [15] and VideoGasNet [48], have shown that background subtraction effectively captures subtle changes in the input and that high-quality masks can be extracted from the resulting foreground using vision–language models (VLMs) together with SAM2 [37]. Therefore, we believe bthat ackground subtraction is suitable for camouflage object segmentation. However, most camouflage object segmentation datasets do not feature a fixed camera; while a moving camera background subtraction method can sometimes work, it may fail under complex camera motion. In addition, if an object does not fully move away to expose the background behind it, a valid background model cannot be built. Although the object’s edges could be shown in the foreground map, camera movements may also highlight these edges, making it hard for the algorithm to distinguish them.
To address such challenges, we turn to another commonly used motion detection method: optical flow. Following Motion Grouping [56] and FlowSAM [52], we employ RAFT [43] to compute optical flow. However, we found that highly repetitive backgrounds can compromise optical flow, thereby diminishing its usefulness. Thus, we combine optical flow with background subtraction, applying the latter (BGS) when there is no camera motion and using optical flow otherwise. To detect camera motion, we used a simple Lucas–Kanade method [26] to track points in the video. The movement was used to estimate affine transformation throughout the video and detect the furthest point the camera reached.
@@ 3.3 Segmentation and Tracking
Given that camouflage can significantly reduce object detection performances for VLM, many frames might result in missed detections. However, previous research has demonstrated that Valiana SAM-2 [41], when guided by explicit prompts, can achieve effective object tracking. We utilized this tracking capability by supplying SAM-2 with all prompts obtained from VLM detections and allowed it to propagate these prompts across all video frames. These prompts consist of the bounding boxes generated by the VLM and the center of mass of the intensity map within each bounding box as a point prompt.
Since forward propagation alone limits object tracking to frames following the initial successful detection, we implemented a bidirectional propagation approach. We provided prompts for both the original forward-playing video and its reversed sequence. Masks generated from both directions were combined using an OR operation, producing the final robust masks across the entire video sequence.

### v2

# ZS-VCOS: Zero-Shot Video Camouflaged Object Segmentation By Optical Flow and Open Vocabulary Object Detection

@@ Abstract
Camouflaged object segmentation presents unique challenges compared to traditional segmentation tasks, primarily due to the high similarity in patterns and colors between camouflaged objects and their backgrounds. Effective solutions to this problem have significant implications in critical areas such as pest control, defect detection, and lesion segmentation in medical imaging. Prior research has predominantly emphasized supervised or unsupervised pre-training methods, leaving zero-shot approaches significantly underdeveloped. Existing zero-shot techniques commonly utilize the Segment Anything Model (SAM) in automatic mode or rely on vision-language models to generate cues for segmentation; however, their performances remain unsatisfactory, due to the similarity of the camouflaged object and the background. This work studies how to avoid training by integrating large pre-trained models like SAM-2 and Owl-v2 with temporal information into a modular pipeline. Evaluated on the MoCA-Mask dataset, our approach achieves outstanding performance improvements, significantly outperforming existing zero-shot methods by raising the F-measure ($F_{\beta}^{w}$) from 0.296 to 0.628. Our approach also surpasses supervised methods, increasing the F-measure from 0.476 to 0.628. Additionally, evaluation on the MoCA-Filter dataset demonstrates an increase in the success rate from 0.628 to 0.697 when compared with FlowSAM, a supervised transfer method. A thorough ablation study further validates the individual contributions of each component. Besides our main contributions, we also highlight inconsistencies in previous work regarding metrics and settings.  [^1] Code can be found on GitHub after publication.
@@ 1 Introduction
Figure 1: Evolution of $F_{\beta}^{w}$ scores over time for supervised and zero-shot methods on an animal dataset. The $F_{\beta}^{w}$ metric is selected for its representativeness and consistent use across all comparison methods. Most zero-shot approaches utilize prior knowledge by explicitly instructing models to detect animals, except for CVP and SAM-2L Auto. Our zero-shot method notably surpasses all previous zero-shot approaches and even outperforms supervised methods, achieving performance close to the test-time semi-supervised upper bound.
Camouflaged object detection and segmentation (COD and COS) is an image detection/segmentation task for objects that are concealed in the background (See Figure 3 for example). It poses significant challenges beyond traditional object detection and segmentation tasks. This increased difficulty primarily stems from the inherent nature of camouflaged objects, which are visually similar to their backgrounds in terms of patterns, colors, and textures [9, 51], effectively blending into their surroundings and complicating accurate identification and delineation. Despite these challenges, effective solutions for COD and COS have considerable real-world significance, especially in critical fields such as defect detection [20], pest control [39], and medical imaging for lesion segmentation [13].
@@
However, camouflage-related tasks in both static and dynamic contexts remain relatively underdeveloped compared to traditional detection and segmentation methods. Most prior work in this area has focused on supervised learning, relying on complex architectures and labelled data. Yet, even these supervised models often struggle with camouflaged objects due to the lack of distinct features. On the other hand, zero-shot methods, which avoid training by using large pre-trained models like SAM and vision-language models, are severely less explored and currently perform worse than supervised methods.
To address this gap, we propose a method that integrates optical flow, a vision-language model, and SAM in a modular pipeline. Each stage of the pipeline uses the output of the previous one to refine its segmentation cues. Rather than relying on any training or fine-tuning, our approach operates in a zero-shot setting and achieves strong performance. On the MoCA-Mask dataset, our method improves mIoU from 0.273 (baseline zero-shot methods) to 0.561. It also outperforms multiple supervised methods, which typically score around 0.422. Furthermore, on the MoCA-Filtered dataset, our method raises the detection success rate from 0.628 to 0.697. These gains highlight the effectiveness of combining motion-based cues with strong foundation models.
In summary, our technical contributions are (1) a zero-shot framework for camouflaged object segmentation in video that surpasses supervised baselines, (2) extensive experimentation on different components and prompting strategies of our methods, and (3) insights demonstrating that properly designed zero-shot pipelines can not only compete with but in some cases outperform traditional supervised approaches.
Additionally, we noticed that previous works often failed to systematically compare their results against other methods evaluated under the same settings (test-time supervised, also known as tracking, and test-time unsupervised). Furthermore, metric calculation in these benchmarks frequently suffered from inconsistent aggregation methods and inadequate handling of special cases. In this work, we highlight these issues, re-evaluate the state-of-the-art methods using a consistent and corrected metric, and ensure a direct and fair comparison between our method and the current state-of-the-art. We urge the research community to adopt standardized evaluation practices to enable clearer and more meaningful comparisons in future studies.
Figure 2: Overview of Our Method.
Figure 3: Visual Comparison Of Our Methods and Previous Supervised Methods.
@@ 2.1 Optical Flow
Optical flow is a technique used to measure pixel movement in videos. It has been used in video processing or recognition for a long time; one of the most significant works is the two-stream network published in 2014 [41]. There are two types of optical flow: sparse and dense optical flow. Sparse optical flow gives a movement vector for points of interest in the image, whereas dense optical flow estimates movement for all pixels in the image. One of the most famous sparse optical flows is the Lucas–Kanade method [27], which uses the assumption that local pixels have similar motion. It can be used in camera motion estimation for panoramic image generation or motion compensation. Dense optical methods, like RAFT [44] and GMFlow [54], can provide movement information for every pixel in the frame, and it has demonstrated promising performance in camouflaged object detection based on movement differences between foreground and background, although their methods relay on training and the perfomance can be furture improved.
[…]
SLT-Net [9] CVPR 22 SV Tr 0.656 0.357 0.021 0.785 0.397 0.310
ZoomNeXt [36] TPAMI 24 SV Tr 0.734 0.476 0.010 0.736 0.497 0.422
TSP-SAM(M+B) [17] CVPR 24 SV Tr 0.689 0.444 0.008 0.808 0.458 0.388
Gao et. al [14] arXiv 25 SV Tr 0.706 0.455 0.011 - 0.495 0.404
SAM2 Tracking [42] arXiv 24 SV Te 0.804 0.691 0.004 - - -
SAM-PM [30] CVPRW 24 SV Tr+Te 0.728 0.567 0.009 0.813 0.594 0.502
Finetuned SAM2-T + Prompts [59] arXiv 24 SV Tr+Te 0.832 0.726 0.005 0.908 0.756 0.652
CVP [43] ACM MM 24 ZS 0.569 0.196 0.031 - - -
SAM-2-S Auto [59] arXiv 24 ZS 0.497 0.201 0.141 0.608 0.202 0.174
LLaVA + SAM2-L [59] arXiv 24 ZS w/ PK 0.624 0.315 0.046 0.688 0.334 0.291
Shikra + SAM2-L [59] arXiv 24 ZS w/ PK 0.502 0.146 0.107 0.590 0.157 0.124
Ours - ZS w/ PK 0.776 0.628 0.008 0.878 0.648 0.550
Table 1: Performance comparison on the MoCA-Mask dataset [9]. “SV Tr” denotes supervised training. “SV Te” denotes supervised testing, where one frame from the video was provided to the model along with prompts. “ZS” indicates zero-shot learning, while “ZS w/ PK” means zero-shot with prior knowledge (since the model already knows it is looking for animals). The grouping of methods is based on settings. Metrics shown in gray represent results from prior work that may contain methodological inconsistencies. These are included for transparency and completeness but should be interpreted with caution. Our method significantly outperforms all zero-shot and even supervisely trained and unsupervisely tested methods.
@@ 2.2.1 Methods Leveraging Optical Flow as Motion Cues
Brox et al. [4] utilized long-term optical flow trajectories combined with clustering for segmenting videos. Ochs et al. [34] applied flow-based motion cues to resolve ambiguities in color-based segmentation. Xiao et al. [52] employed optical flow cues to reinforce target frame representations. Yang et al. [58] used both optical flow and RGB input to assist video object segmentation. FlowI-SAM and FlowP-SAM [53] utilized optical flow either exclusively as input (FlowI-SAM) or as a prompt guiding segmentation of RGB frames (FlowP-SAM). Both FlowI-SAM and FlowP-SAM handled standard and camouflaged objects effectively.
@@ 2.4 Camouflage Object Detection and Segmentation
Unlike regular object detection and segmentation, camouflage object tasks are significantly more challenging because the foreground usually seamlessly blends into the background. There are two tasks in camouflage object detection: image-based camouflage object detection (usually referred to as COD) and video-based camouflage object detection (VCOD). They could also be extended to segmentation, namely COS and VCOS. VCOD/S allows the model to use motion cues to detect the foreground but also brings in the challenges of temporal changes [51].
@@ 2.4.1 Datasets
There are two major datasets and 3 variants in video camouflage object detection: Camouflaged Animal Dataset (CAD) [3] and Moving Camouflaged Animal Dataset (MoCA) [22]. However, MoCA is an object detection dataset but not a segmentation dataset. It contains some non-camouflaged animals or animals that do not have locomotion. Thus, two variances of MoCA were proposed. MoCA-filtered [57] and MoCA-Mask [9].
MoCA-filtered mainly removed non-locomotive videos from the dataset, with additional processing such as cropping away logos and borders, resampling frames, and incorporating the bounding boxes. Since it still lacks segmentation masks, papers using this dataset ([57] [53]) used the detection success rate based on the IoU threshold to evaluate the results. MoCA-Mask improved MoCA by removing scenes with obvious animals and converting bounding boxes into masks. In addition to ground truth masks provided every 5 frames, they also used bidirectional optical flow to generate pseudo masks for unlabelled frames.
@@ 2.4.2 Algorithms
SLT-Net [9] is a supervisely trained and unsupervisely tested model. It argued that when using optical flow and homography, the error might be accumulated from both the motion estimation and segmentation. Thus, they proposed to use a unified framework for both motion estimation and segmentation. Additionally, they used a long-term spatiotemporal transformer to refine short-term predictions, although this long-term module provides marginal improvement. ZoomNeXt [36] is an improved version of ZoomNet [35], mainly adapted from image-based COS to video-based COS and improved performance by introducing more structural extensions. ZoomNeXt is trained on both image COS datasets and video COS datasets, including MoCA-Mask [9]. ZoomNet and ZoomNeXt both use zooming to capture features at different scales. Similar to SLT-Net, they are both supervisely trained and unsupervisely tested methods.
Previous studies have sometimes failed to clearly differentiate between test-time semi-supervised and unsupervised tasks, despite their differing levels of difficulty. For example, SAM-PM [30], requiring supervision during both training and inference, reported state-of-the-art results compared with SLT-Net [9]. However, SLT-Net operates under supervised training but unsupervised testing conditions. This fundamental difference in evaluation criteria renders direct comparisons between these two methods somewhat inequitable. Although the authors of SAM-PM described their method as semi-supervised (which we refer to in this paper as test-time supervised), they did not clearly acknowledge this distinction when making comparisons or drawing conclusions.
@@
For zero-shot unsupervised testing neither training nor inference is supervised), Chain of Vision Perception (CVP) [43] represents an early effort employing vision-language models (VLMs) for COD/S tasks, with a primary focus on images rather than videos. CVP prompts a vision-language model to identify the location of camouflaged objects. Subsequently, these locations are refined and given to a segmentation model. Properly designed prompting can further enhance the model’s performance. CVP achieved higher performance than several supervised methods on datasets such as CAMO [23], COD10K [12], and NC4K [28]. However, its results on the MoCA-Mask were suboptimal, with a weighted F-score ($F_{\beta}^{w}$) of 0.196. Zhou et al. [59] improved upon this by employing LLaVA [25] or Shikra [5] as the vision-language model and utilizing SAM-2 for segmentation. A similar approach is evident in Grounded SAM [26], which integrates Grounding DINO and SAM for open-vocabulary segmentation of regular objects.
@@
Motion Detection Mean Subtraction Momentum Tracking $S_{\alpha}\uparrow$ $E_{\phi}\uparrow$ mIoU $\uparrow$
@@ 3.1 Motion Detection
Our method builds upon LangGas [16]. Gas leakage shares many similarities with a camouflage object: they both have low contrast against the background, but they often have different relative motion with respect to the background. Previous studies, including LangGas [16] and VideoGasNet [49], have shown that background subtraction effectively captures subtle changes in the input. High-quality masks can then be extracted from the resulting foreground using vision–language models (VLMs) together with SAM2 [38] [16]. However, traditional BGS methods can only be used in fixed camera settings, and most camouflage object segmentation datasets and real-world applications do not feature a fixed camera; while a moving camera background subtraction method can sometimes work, it may fail under complex camera motion. In addition, if an object does not fully move away to expose the background behind it, a valid background model cannot be built.  [^2] Although the object’s edges could be shown in the foreground map, camera movements may also highlight these edges, making it hard for the algorithm to distinguish them.
To address such challenges, we turn to another commonly used motion detection method: optical flow. By tracking the movement of each pixel between two adjacent frames, optical flow can show different movement patterns in the image. Following Motion Grouping [57] and FlowSAM [53], we employ RAFT [44] to compute optical flow. However, we found that highly repetitive backgrounds or videos with margins can compromise RAFT optical flow, thereby diminishing its usefulness. Thus, we combine optical flow with background subtraction, applying the latter (BGS) when there is no camera motion and using optical flow otherwise. To detect camera motion, we use a simple Lucas–Kanade method [27] to track points in the video. The movement is used to estimate the affine transformation throughout the video and detect the furthest point the camera reached.
@@ 3.3 Segmentation and Tracking
Given that camouflage can significantly reduce object detection performance for VLM, many frames might result in missed detections. However, previous research [59] [30] [42] has demonstrated that vanilla SAM-2 [42], when guided by explicit prompts, can achieve effective object tracking. We utilized this tracking capability by supplying SAM-2 with all prompts obtained from VLM detections and allowed it to propagate these prompts across all video frames. These prompts consist of the bounding boxes generated by the VLM and the center of mass of the intensity map within each bounding box as a point prompt.
Since forward propagation alone limits object tracking to frames following the initial successful detection, we implemented a bidirectional propagation approach. We provided prompts for both the original forward-playing video and its reversed sequence. Masks generated from both directions are combined using an OR operation, producing the final robust masks across the entire video sequence.